OpenStudy (anonymous):

A falling skydiver has a mass of 105 kg. What is the magnitude of the skydiver's acceleration when the upward force of air resistance has a magnitude that is equal to one-half of his weight?

OpenStudy (anonymous):

Pop in and pop out.. F_free = m * g - F_air = m*a F_air = 1/2*m a = (105kg * 9.8 m/s^2 - .5*105)/105kg a = 9.3 m/s^

OpenStudy (anonymous):

^___^

OpenStudy (anonymous):

hey.. mark me up with a metal.. I've been trying to pick them up today.. "Good Answer" thing.. :D

OpenStudy (anonymous):

i got you :D

OpenStudy (anonymous):

swweeeeettt.. :DD

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

See me roll up.. on my Segway..... *Guess I'm just too white and nerdy*

OpenStudy (anonymous):

Hahahaha!

OpenStudy (anonymous):

I'm actually listening to that song now.. :P

OpenStudy (anonymous):

hey i dont quite understand it why is it that a = (105kg * 9.8 m/s^2 - .5*105)/105kg and not a = (mg - F_air)/105 = (mg - 0.5*mg)/105 = (105*9.8 - 0.5*105*9.8)/105 = 4.9 ms^-2

OpenStudy (anonymous):

since F_air = 1/2 of his weight?

OpenStudy (anonymous):

F_free = m * g - F_air = m*a F_air = 1/2*m So, for the first equation, you substitute the F_air to get the same equation just with the F_air substituted: F_Free=m*g-(1/2*m) =m*a Hope this helps!